OpenStudy (anonymous):

Solve the equation: 29+x/72+x=.5

OpenStudy (zarkon):

14

OpenStudy (anonymous):

If the Tigers won 2 out of every 3 (or 2/3) of their remaining games, how many games would it take them to reach a winning percentage of .500? (EQUATION)

OpenStudy (zarkon):

42

OpenStudy (zarkon):

42*(2/3)=28 (29+28)/(72+42)=.5